Jennifer deposits 15.3. 6148 € in an account with an annual interest rate of 1.76 %. At the end of that year, Jennifer withdraws her deposit (including the capital and acrrued interest). Her friend Janet wants to withdraw from her own account at the same time exatcly the same amount of money (including the capital and acrrued interest) as Jennifer. Janet makes her deposit 20.7. and the interest rate of her account is 2.1 % annually. The tax at source rate is 30 %. What is the amount of Janet's deposit?


1
Expert's answer
2022-02-07T15:31:48-0500

1. Jannifer deposit at the end of year: 6148 + (6148*1.76%)*(100%-30%) = 6223,74 €

2. Make the equation of Janet deposit: x + (x*2.1%)*(100%-30%) = 6223.74

3. x + 0.0147*x = 6223.74

4. 1.0147*x = 6223.74

5. x = 6223.75/1.0147 => x = 6133.58 €


Answer: The amount of Janet's deposit is 6133.58 €